# Reputation in the Shadow of Exit

### v5

@@ 3.2 Guaranteeing Stackelberg Payoffs
Suppose $v(a,c)=v(c)$, $L(a,c)=\mathcal{L}(c)$ for some $\mathcal{L}(c)$, $|\mathcal{N}(0)|=1$, and $V^{*}=V_{*}$. Then the following are equivalent.
@@
    There exists $\overline{\mu}_{0}>0$ such that for all $\mu_{0}<\overline{\mu}_{0}$, $\delta>\overline{\delta}$, $\underline{U}^{M}=V^{*}$.
@@
The proof simply verifies that, under the stated conditions, the incentive compatibility conditions of Proposition 1 and Theorem 1 coincide. It is delegated to the Proof of Corollary 1
@@
3.3 Non-Markov equilibrium
Theorem 1 focuses on Markov strategies and leaves open the question of when reputation can help refine payoffs away from the Markov case. Without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markov strategies? I next show the answer is no—even with $c^{*}$ incentive compatibility, one cannot refine the payoff set in general, as the long-run player’s payoff may be arbitrarily low.
Proposition 3 (Anti-Commitment.).
Fix payoffs $(u,v)$ such that there exists $c$ where
$\min_{a\in\mathcal{N}(c)}v(a,c)>0$. Then for any $\varepsilon>0,M>0$, there exists $f$ such that
  1. (1)
    The game admits a folk theorem and $V^{*}\geq M$.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq\varepsilon$.
The first assumption is a nontriviality condition guaranteeing the existence of a non-Nash commitment posture, conditional on a high probability of survival. The limit in the second statement guarantees the probability of commitment is not “so high” that what the rational long-run player does will not affect their payoffs (i.e. if $\mu_{0}(\omega_{c^{*}})\approx 1$).
Proposition 3 is stronger than the simple statement that the long-run player’s payoff can be arbitrarily low, which can be achieved by simply setting $f(a,c)\leq\eta$ for all pairs $(a,c)$ and arbitrarily low $\eta$. In particular, the first statement guarantees that the long-run player attains a low payoff even if (1) the folk theorem holds, so they can attain their Stackelberg payoff, and (2) such a Stackelberg payoff is arbitrarily high. In this way, Proposition 3 guarantees that in sequential equilibrium, even in the prescence of reputational incentives, the payoff set can become large absent strong assumptions (stronger than $c^{*}$ incentive compatibility) on the exit function. One consequence of Proposition 3 is that as $\mu_{0}(\omega_{R})\to 1$, the set of sequential equilibria payoffs with reputation converges to the set of sequential equilibria payoffs without, in stark contrast to the baseline payoff selection results of Fudenberg and Levine (1989) without exit.
Proposition 3 highlights the two important roles of Markov equilibrium in establishing the positive reputation result of Theorem 1. First, it disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward tomorrow actions which are not the commitment action. Second, it ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period (c.f. the discussion succeeding Theorem 1. Markov beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today. The proof is relegated to the PROOF OF PROPOSITION 3
@@ 4 The Chain Store Game
Formally, suppose in each period the long-run player ($I$) chooses to fight or accomodate an entrant, but only after the entrant ($E$) has decided whether or not to enter. Let stage game payoffs be as follows, where the first entry is the long-run player’s payoff. [^12] These specific payoffs for the chain store game are due to Harry Pei’s Lecture Notes. Interpret here “accomodate” as the action 0.
@@
Moreover, define $f(a,c)$ to take the values in 1, where $\varphi,\varepsilon,\eta$ are strictly positive and $1-\eta\geq\max\{\varphi,\varepsilon\}$, so survival occurs with higher probability if the entrant chooses Out.
@@
Table 1: Exit Probabilities for the Chain Store Game
Note that these payoffs imply the long-run player’s commitment action $c^{*}$ is $F$ for all $\delta$, but that accomodate is the unique Nash equilibrium. These observations (in addition to ranking $F\succ A=0$) imply that Assumption 1 is satisfied. The chain store game admits the following result.
@@
Moreover, there exists $e(\cdot)$, vanishing at 0 such that
$\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq e(\varepsilon).$
@@
Three observations are in order. First is that in Proposition 4, the key parameter was the probability of exit given that the incumbent entered the game. Intuitively, varying $\varepsilon$ varies the effective discount rate but also the incentives to take different actions; lower $\varepsilon$ increase the “cost” of Nash equilibrium and make reputation formation more attractive to the long-run player, especially relative to fighting. Mathematically, this is evidenced through the likelihood ratio $\frac{\varphi}{\varepsilon}$; as $\varepsilon\to 0$, the likelihood ratio of survival after fighting explodes relative to that after accommodating. Second is the contrast of Proposition 4 with known results from reputation a la Fudenberg and Levine (1989). In particular, the second part of the proposition highlights how reputation can fail with exit even with extremely patient players and high likelihood ratios for survival conditional on the commitment action. This shows that the Markovian refinement is critical for any hope of selecting a unique equilibrium, exactly because it rules out nonstationary equilibrium where the exit cost becomes “sunk” to the long-run player, who chooses their stage-game action conditional on the fact they have already survived today. Finally, I am able to give a unique behavioral prediction in Markov strategies in this game in which the incumbent always fights and short-run players never enter (i.e. deterrence is successful), unlike the indeterminancy of behavioral predictions in reputation games in general (c.f. Pei and Li (2021)).
@@ 5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\overline{c}]$ and moreover suppose $\kappa(\overline{c})>1$, so that the government never chooses the maximally feasible value of punishment. [^13] Note $C=[0,\overline{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(0)$, after which it is sufficient to assume $C=\{0,c^{*}(0)\}$ for the remainder of the application. [^14] This is an Inada-type condition that ensures the government’s problem is always well-defined
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(\theta)=\tau^{2}$.  [^15] In practice, I will consider the case where $\tau^{2}$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfect signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
@@
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals (i.e. player $i$’s thought about others’ signals after receiving $x$). The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the state is not in a dominance region.
Assumption 2 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable as the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature. [^16] I am not aware of a global games setting studied in the literature that does not satisfy these assumptions; for example it nests normal-normal information structures.
Having seen their private signals, the dissidents independently and simultaneously choose an action $a_{i}\in\{0,1\}$ in each period, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents defect (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt is successful, is given by $f(A,c)=\mathbb{P}_{\theta}(A\geq\theta)$, and is independent of $c$.
@@ 5.2 Selecting Equilibria
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(0)$. I also assume by fiat that $c^{*}\neq 0$, which is assumed throughout the main exposition, so that the commitment type is distinguishable from one playing static Nash. [^17] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$.
@@
The result stands in contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the “survival dominant” equilibrium (i.e. survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markov equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculator’s sense the regime is “likely weak.”
@@ 6 Discussion
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Understanding the set of Nash equilibria in this game and their payoff bounds is potentially promising. Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the repeated trust game, my findings can give predictions about when the long-run player can teach short-run players to trust them even when a lack of trust might prematurely end the game. Alternatively, Theorem 1 can be used to give conditions under which Markov equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markov predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
@@
By construction short-run players are myopically best replying. Clearly at histories where $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ the long-run player’s strategy is sequentially rational. At a history where only $c^{*}$ has been played, Assumption 1 implies the long-run player’s most profitable deviation is to play 0; this is unprofitable so long as
@@
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $\underline{V}_{*}$. Note this argument holds even for the discounted repeated game, for any discount rate. ∎
@@
First, note that if $\mu_{0}(\omega_{R})=1$, then it must be that in every sequential equilibrium $\mu(h^{t})(\omega_{R})=1$ as well (as the agent has a degenerate prior belief). Now fix any history $h^{t}$; at any successor history $h^{s},h^{s^{\prime}}\succ h^{t}$, it must be that $\sigma(h^{s})=\sigma(h^{s^{\prime}})$, so in particular future play cannot vary based on play today. Thus, the long-run player’s action must be myopically optimal, and hence $\sigma_{L}(h^{t})=0$ at any history $h^{s}$. Short-run players must play the same myopic best response at each history, so in equilibrium $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(0)$ always. Hence all equilibria are repeated static Nash. Public randomization convexifies the payoff set. ∎
@@
Let $(\sigma,\mu)$ be Markov perfect. By Proposition 2, any continuation strategy at histories $h^{t}$ where $\mu(h^{t})(\omega_{R})$ must be a repeated static Nash profile, so in particular $\sigma_{L}(h^{t})=0$.
Consider now a history $h^{t}$ where $\mu(h^{t})(\omega_{R})<1$ (noting these are all on path), and suppose the lemma is false. Then there exists some on-path continuation history $h^{t}$ where at $h^{t}\succ h^{t-1}$, $h^{t-1}=\{c^{*}\}_{t=0}^{t-2}$ but $h_{t}=c\not\in\{0,c^{*}\}$. Thus, $\mu(h^{t})(\omega_{R})=1$ and the long-run player’s continuation value at $h^{t}$ must be the same as their continuation value at $\{h^{t-1},0\}$. But then playing 0 instead of $c$ for sure is a strictly profitable deviation by Assumption 1, contradicting sequential rationality of $(\sigma,\mu)$. ∎
@@
$\mathcal{T}=\{t:h^{t}\text{ is clean and }\sigma_{L}(h^{t})\text{ mixes.}\}$
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\overline{V}_{0}$ hereafter, since they have revealed their rationality) and taking $c^{*}$ (and obtaining some other future continuation value). In particular, for $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$, we have
@@
Yet we also know from Proposition 2 that
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]\leq f(a,0)[v(a,0)+\overline{V}_{0}]$
and from our observation above that
@@
Yet the assumption the long-run player satisfies $c^{*}$ IC implies
@@
for $a$ since $a\in\mathcal{N}(\alpha c)$. Taking everything together and using $c^{*}$-incentive compatibility implies for all $N$ sufficiently large that
@@
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}$ when they play 0. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\alpha c)$ for some $\alpha\in[0,1]$, we have
$f(a,0)[v(a,0)+\underline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V^{*}]$
This again contradicts $c^{*}$ IC. Hence the long-run player never mixes. ∎
@@
Suppose the condition fails. There are three cases. First, suppose the inequality fails strictly at $\alpha=0$. Then in particular
@@
for $a\in\mathcal{N}(0)$. This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\overline{V}_{0}$ afterwards. In particular, it implies the following is a Markov equilibrium as $\mu_{0}(\omega_{R})\to 1$:

### v6

@@ 3.3 Guaranteeing Stackelberg Payoffs
Suppose $v(a,c)=v(c)$, $L(a,c)=\mathcal{L}(c)$ for some $\mathcal{L}(c)$, $|\mathcal{N}(0)|=1$, and $V^{*}=V_{*}$. Then the following are generically equivalent:
@@
    There exists $\overline{\mu}_{0}>0$ such that for all $\mu_{0}<\overline{\mu}_{0}$, $\underline{U}^{M}=V^{*}$.
@@
The proof simply verifies that, under the stated conditions, the incentive compatibility conditions of Proposition 1 and Theorem 1 coincide modulo a weak equality when the enforceability constraint exactly binds in Proposition 1. But the enforceability constraint binding is nongeneric, and hence Corollary 2 holds so long as primitives are not such that $V^{*}L(a^{*},c^{*})-\underline{V}_{0}=v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*})$. The formal proof is delegated to the Proof of Corollary 2
@@ 4 The Chain Store Game
Formally, suppose in each period the long-run player ($I$) chooses to fight or accomodate an entrant, but only after the entrant ($E$) has decided whether or not to enter. Let stage game payoffs be as follows, where the first entry is the long-run player’s payoff [^17] These specific payoffs for the chain store game are due to Harry Pei’s Lecture Notes.. Interpret here “accomodate” as the action 0181818The chain-store game is a sequential move game; the analysis applies to the normal form of the extensive form game..
@@
Moreover, define $f(a,c)$ to take the values in Table 1, where $\varphi,\varepsilon,\eta$ are strictly positive and $1-\eta\geq\max\{\varphi,\varepsilon\}$, so survival occurs with higher probability if the entrant chooses Out.
@@
Table 1: Survival Probabilities for the Chain Store Game
Note that these payoffs imply the long-run player’s commitment action $c^{*}$ is $F$ for all $\delta$, but that Accomodate is dominant. These observations (in addition to ranking $F\succ A=0$) imply that Assumption 1 is satisfied. The chain store game admits the following result.
@@
Moreover, there exists $e(\varphi,\varepsilon)$, vanishing at $(0,0)$ such that
$\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq e(\varphi,\varepsilon).$
@@
Three observations are in order. First is that in Proposition 4, the key parameter was the probability of exit given that the incumbent entered the game. Intuitively, varying $\varepsilon$ varies the effective discount rate but also the incentives to take different actions; lowering $\varepsilon$ increases the “cost” of Nash equilibrium and makes reputation formation more attractive to the long-run player, especially relative to fighting. Mathematically, this is evidenced through the likelihood ratio $\frac{\varphi}{\varepsilon}$; as $\varepsilon\to 0$, the likelihood ratio of survival after fighting explodes relative to that after accommodating. Second is the contrast of Proposition 4 with known results from reputation a la Fudenberg and Levine (1989). In particular, the second part of the proposition highlights how reputation can fail with exit even with extremely patient players and high likelihood ratios for survival conditional on the commitment action. This shows that the Markovian refinement is critical for any hope of selecting a unique equilibrium, exactly because it rules out nonstationary equilibrium where the exit cost becomes “sunk” to the long-run player, who chooses their stage-game action conditional on the fact they have already survived today. Finally, I am able to give a unique behavioral prediction in Markov strategies in this game in which the incumbent always fights and short-run players never enter (i.e. deterrence is successful), unlike the indeterminancy of behavioral predictions in reputation games in general (c.f. Pei and Li (2021)).
@@ 5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\overline{c}]$ and moreover suppose $\kappa(\overline{c})>1$, so that the government never chooses the maximally feasible value of punishment [^19] Note $C=[0,\overline{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(0)$, after which it is sufficient to assume $C=\{0,c^{*}(0)\}$ for the remainder of the application. [^20] This is an Inada-type condition that ensures the government’s problem is always well-defined.
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(\theta)=\tau$.  [^21] In practice, I will consider the case where $\tau$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfect signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
@@
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals (i.e. player $i$’s thought about others’ signals after receiving $x$). The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the average state is not in a dominance region.
Assumption 2 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable as the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature [^22] I am not aware of a global games setting studied in the literature that does not satisfy these assumptions; for example it nests normal-normal information structures..
Having seen their private signals, the dissidents independently and simultaneously choose an action $a_{i}\in\{0,1\}$ in each period, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents attack (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt fails, is given by $f(A,c)=\mathbb{P}_{\theta}(A\leq\theta)$, and is independent of $c$.
@@ 5.2 Selecting Equilibria
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(0)$. I also assume by fiat that $c^{*}\neq 0$, which is assumed throughout the main exposition, so that the commitment type is distinguishable from one playing static Nash [^23] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$..
@@
The result stands in contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the “survival dominant” equilibrium (i.e. survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markov equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculators sense the regime is “likely weak.”
@@ 6 Discussion
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the repeated trust game, my findings can give predictions about when the long-run player can teach short-run players to trust them even when a lack of trust might prematurely end the game. Alternatively, Theorem 1 can be used to give conditions under which Markov equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markov predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
@@
By construction short-run players are myopically best replying. Clearly at histories where $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ the long-run player’s strategy is sequentially rational. At a history where only $c^{*}$ has been played, Assumption 1 implies the long-run player’s most profitable deviation is to play 0 (since continuation play is the same under any deviation); this is unprofitable so long as
@@
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $V^{*}$. This finishes the proof. ∎
@@
First, note that if $\mu_{0}(\omega_{R})=1$, then it must be that in every sequential equilibrium $\mu(h^{t})(\omega_{R})=1$ as well for all on-path histories (as the agent has a degenerate prior belief). Now fix any history $h^{t}$; at any successor history $h^{s},h^{s^{\prime}}\succ h^{t}$, it must be that $\sigma(h^{s})=\sigma(h^{s^{\prime}})$, so in particular future play cannot vary based on play today. Thus, the long-run player’s action must be myopically optimal, and hence $\sigma_{L}(h^{t})=0$ at any history $h^{s}$. Short-run players must play the same myopic best response at each history, so in equilibrium $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(0)$ always. Hence all equilibria are repeated static Nash. Public randomization convexifies the payoff set. ∎
PROOF OF PROPOSITION 3
Proof.
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be some best reply to it where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Define an exit probability function $f$ as follows: $f(a^{*},c^{*})=1-\eta$, $f(a^{*},0)=\eta$, and $f(\underline{a})=\eta$ for every $\underline{a}\in\mathcal{N}(0)$. Let the other values of $f$ be chosen arbitrarily. Note for $\eta$ sufficiently small, playing $c^{*}$ must be preferred to playing 0 since
$\frac{v(a^{*},c^{*})(1-\eta)}{1-(1-\eta)}\geq\frac{v(\underline{a},0)\eta}{1-\eta}.$
This in particular implies for all $\eta>0$, there is an optimal commitment posture $c^{*}\neq 0$ and in particular one which guarantees a payoff of at least $v(a^{*},c^{*})>0$ in each period conditional on surviving.
$\mathcal{L}(a^{*},c^{*}(\eta))\left(v(a^{*},c^{*}(\eta))+V^{*}\right)\geq v(a^{*},0)+\underline{V}_{0}(\eta)$
The right hand side is bounded from above in $\eta$ while the left hand side diverges at $\eta\to 0$, since it is equivalent to
$\frac{1-\eta}{\eta}(v(a^{*},c^{*}(\eta))+V^{*}).$
Hence under this survival probability function $f^{*}$, $c^{*}(\eta)$ is enforceable for all $\eta$ sufficiently small.
  1. (1)
    Have the long-run player play 0 in the first period.
  2. (2)
    Have the short-run players play $a^{*}(\mu_{0}(\omega_{c^{*}(\eta)})c^{*}(\eta)))$, i.e. best respond to their prior belief supposing the rational type plays 0.
  3. (3)
    Regardless of the long-run player’s action, given them a payoff of $V_{*}$ tomorrow.
Clearly $V_{*}$ is sustainable by the folk theorem tomorrow. Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action $\underline{a}$ in $\mathcal{N}(0)$. Moreover, the long-run player will play 0 because $f(\underline{a},c)$ is constant in $c$ and their continuation payoff is independent of their choice of action, so they will play their dominant action 0 today, i.e.
$\{0\}=\operatorname*{\arg\!\max}_{c\in C}f(\underline{a})[v(\underline{a},c)+V_{*}]$
by the assumption $0\in\operatorname*{\arg\!\max}_{c\in C}v(a,c)$ at $\underline{a}$. Thus the long-run player gets a payoff of exactly
which diverges as $\varsigma\to 0$. ∎
@@
Let $(\sigma,\mu)$ be Markov perfect. By Proposition 2, any continuation strategy at histories $h^{t}$ where $\mu(h^{t})(\omega_{R})=1$ must be a repeated static Nash profile, so in particular $\sigma_{L}(h^{t})=0$.
Consider now a history $h^{t}$ where $\mu(h^{t})(\omega_{R})<1$ (noting these are all on path), and suppose the lemma is false. Then there exists some on-path continuation history $h^{t}$ where at $h^{t}\succ h^{t-1}$, $h^{t-1}=\{c^{*}\}_{t=0}^{t-2}$ but $h_{t}=c\not\in\{0,c^{*}\}$. Thus, $\mu(h^{t})(\omega_{R})=1$ and the long-run player’s continuation value at $h^{t}$ must be the same as their continuation value at $\{h^{t-1},0\}$. But then playing 0 instead of $c$ for sure is a strictly profitable deviation by Assumption 1 (it increases the long-run player’s stage game payoff without affecting their continuation value), contradicting sequential rationality of $(\sigma,\mu)$. ∎
@@
[…]
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c^{*}$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\overline{V}_{0}$ hereafter, since they have revealed their rationality) and taking $c^{*}$ (and obtaining some other future continuation value). In particular, for $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$, we have
@@
Yet we also know from Proposition 2 that for all $\varepsilon>0$, there exists $N$ sufficiently large such that
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t_{n}},0\}|\sigma)]\leq f(a,0)[v(a,0)+\overline{V}_{0}]+\varepsilon$
and (combined with from our observation above) that
@@
Yet the assumption the long-run player satisfies $c^{*}$ incentive compatibility implies
@@
for $a$ since $a\in\mathcal{N}(\beta c)$. Taking everything together and using $c^{*}$-incentive compatibility implies for all $N$ sufficiently large and $\varepsilon$ sufficiently small
@@
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}$ when they play $c^{*}$. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\beta c)$ for some $\beta\in[0,1]$, we have
$f(a,0)[v(a,0)+\bar{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V_{*}]$
This again contradicts $c^{*}$ incentive compatibility. Hence the long-run player never mixes. ∎
@@
Suppose the condition fails. There are three cases. First, suppose the inequality fails strictly at $\beta=0$. Then in particular
@@
for $a\in\mathcal{N}(0)$. This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\overline{V}_{0}$ afterwards. In particular, it implies there exists $\bar{\mu}_{0}<1$ such that the following is a Markov equilibrium for all $\mu_{0}(\omega_{R})>\bar{\mu}_{0}$: